6 over x squared minus 8x = 1 over x + 3 over x squared minus 8x
Answer by RAY100(1637)

6/(x^2 -8x) =1/(x+3)/(x^2 -8x)= (x^2 -8x)/(x+3),,,,cross multiply
.
6(x+3) = (x^2 -8x)(x^2-8x)
.
6x +18 = x^4 -16x^3 +64x^2
.
0= x^4 -16x^3 +64x^2 -6x -18
.
solving using graphing calc,,,
.
x=-.32,,,+.59,,,+6.97,,+8.94
.